# TG-NAS: Leveraging Zero-Cost Proxies with Transformer and Graph Convolution Networks for Efficient Neural Architecture Search

### v1

@@ Abstract
Neural architecture search (NAS) is an effective method for discovering new convolutional neural network (CNN) architectures. However, existing approaches often require time-consuming training or intensive sampling and evaluations. Zero-shot NAS aims to create training-free proxies for architecture performance prediction. However, existing proxies have suboptimal performance, and are often outperformed by simple metrics such as model parameter counts or the number of floating-point operations. Besides, existing model-based proxies cannot be generalized to new search spaces with unseen new types of operators without golden accuracy truth. A universally optimal proxy remains elusive. We introduce TG-NAS, a novel model-based universal proxy that leverages a transformer-based operator embedding generator and a graph convolution network (GCN) to predict architecture performance. This approach guides neural architecture search across any given search space without the need of retraining. Distinct from other model-based predictor subroutines, TG-NAS itself acts as a zero-cost (ZC) proxy, guiding architecture search with advantages in terms of data independence, cost-effectiveness, and consistency across diverse search spaces. Our experiments showcase its advantages over existing proxies across various NAS benchmarks, suggesting its potential as a foundational element for efficient architecture search. TG-NAS achieves up to 300$\times$ improvements in search efficiency compared to previous SOTA ZC proxy methods. Notably, it discovers competitive models with 93.75% CIFAR-10 accuracy on the NAS-Bench-201 space and 74.5% ImageNet top-1 accuracy on the DARTS space.
Keywords:
zero-cost proxy neural architecture search transformer
@@ 1 Introduction
Deep convolutional neural networks (CNN) have achieved incredible performance in computer vision, speech recognition, object detection, and many other fields [16, 25, 34, 12]. Since then many manually designed CNN topologies that target either high performance or high computational efficiency have been proposed [10, 38]. With larger and deeper expert-designed CNN models, people can easily achieve state-of-the-art performance in image classification, object detection, and many other applications. However, manual network architecture design is costly and requires a significant amount of time, expertise, and computing resources. It becomes even more challenging if the neural network is designed for hardware platforms with stringent resource and energy constraints. Thus neural architecture search (NAS) was proposed to automate the model design and provide an opportunity to explore specialized model architectures within certain performance and resource constraints [41, 23]. Despite the advancements in automated search techniques, the majority of neural architecture search (NAS) approaches still face the challenge of time-consuming training and evaluation. In the early stage, the prevailing methodology involved iteratively sampling DNN architectures from the search space, training them, and using their performance to guide the search process [22]. These approaches were primarily based on reinforcement learning (RL) or evolutionary algorithms.
@@
In recent years, zero-shot proxies [46, 15, 31, 35] have emerged as a promising technique in neural architecture search (NAS). Zero-shot proxies use lightweight computation, typically a single forward pass with one minibatch of data, to assign a score to each DNN architecture in the search space. The scores obtained from these proxies are expected to correlate with the final accuracy of candidate architectures [46, 15, 31]. This approach not only conserves computational resources but also accelerates the overall NAS process, enabling researchers to explore a larger number of architectures. However, recent general zero-shot proxies struggle to consistently perform well for practical uses. Many of them cannot outperform a naïve proxy like the number of parameters (Params) and are often data-dependent and not fast enough. The generally high dependence and correlation of different zero-cost (ZC) proxies from similar domains, such as gradient information or kernel methods [18, 7], may also compromise the effectiveness of zero-shot NAS approaches.
@@
To address these limitations of zero-cost performance indicators, many model-based performance predictors have emerged. These predictors rely on training models to forecast the final validation accuracy of an architecture solely based on its topology and operators. Common models include Gaussian processes [20], deep neural networks [30, 40], among others. However, these methods are not zero-shot and cannot be considered zero-cost since they typically require hundreds of fully-trained architectures as training data, resulting in prolonged search times. Additionally, model-based prediction has often been used only to accelerate NAS query times, given their lack of generalization and inability to handle unseen operators that are not encoded during predictor model training.
@@
This leads us to deep think: how can we develop a pure universal prediction model capable of predicting any given architecture? Is it possible to develop a generally applicable model-based predictor that can handle the aforementioned unseen operator and break the generalization limitations? In other words, why let model-based predictors settle for a subroutine when it has the potential to take the center stage? Our response to these questions is TG-NAS, a model-based predictor-only NAS framework that can be generalized to new search spaces and unseen operators without retraining. It features a general transformer operator encoder and a graph convolution network (GCN) trainer. This novel approach offers an exceptionally efficient and zero-cost solution to general neural architecture search problems and bridges the gap between model-based predictors and ZC proxies. We will open source this work to facilitate future research. Link to the source code can be found in the supplementary materials.
The main contributions of TG-NAS can be summarized as follows:
@@
  • We propose a universally applicable, data-independent performance predictor model that can handle unseen operators in new search spaces without retraining. The model features a graph convolutional network (GCN), an architecture encoder, and a transformer-based operator embedding generator.
  • Our proposed model acts as a zero-cost proxy, guiding architecture search. This opens up a new space for prediction model-only architecture search, and we provide a comprehensive analysis on proxy independence.
  • We propose a pruning-evolutionary hybrid searching method, enabling quick and accurate identification of the best architecture within the search space.
@@
  • Our analysis of TG-NAS proxies demonstrates remarkable performance on both NAS-Bench-201 and DARTS spaces with CIFAR-10/ImageNet datasets. TG-NAS achieves up to 300×300$\times$300 × faster search compared to other zero-cost proxies and up to 331,200×331,200$\times$331 , 200 × faster than other NAS methods, while maintaining high accuracy.
@@ 2.1 Neural Architecture Search (NAS)
Conventional NAS algorithms are an important step in automating the design of neural architectures for a given task[46]. A typical NAS contains a search strategy that selects the candidate neural architecture from a predefined search space and an estimation strategy that enables the performance estimation for the current candidate.
Search space can be categorized as the following types: macro search space [13], chain-structures search space [38], cell-based search space [28, 53, 8], and hierarchical search space[27]. Among those, cell-based search space is the most popular one in NAS [46]. The searchable cells (a directed acyclic graph of operations) make up the microstructure of the search space while the macrostructure (ways and number of cells to stack together) is fixed. For example, NAS-Bench-101 [53] contains 423,624 unique architectures, and each cell consists of 7 nodes (each node chosen from 3 operators). In NAS-Bench-201[8], there are 15,625 cell candidates and each cell consists of 4 nodes (each node chosen from 5 operators). In contrast, the DARTS search space [28] is more expansive, featuring approximately $10^{18}$ architectures. It consists of two cells, each containing seven nodes. The first two nodes receive inputs from previous layers, and the intermediate four nodes can assume any DAG structure, each having two incident edges. The last node serves as the output node, and each edge can take on one of eight operations. In this work, we perform our experiments on those three search spaces for different purposes.
The search strategy in NAS has been widely explored. There are some well-known black-box optimizations, such as random selection with full training, reinforcement learning, evolution, Bayesian optimization, and Monte Carlo tree search. With these search strategies, we still need to train the searched architecture and use the performance result to guide the search, which is a time-consuming process. To overcome the training time bottleneck, one-shot techniques were introduced as an estimation strategy. These techniques involve training a single (one-shot) supernet, which is an over-parameterized architecture that encompasses all possible architectures in the search space [3, 28, 2]. Once the supernet has been trained, each architecture in the search space can be evaluated by inheriting its weights from sampling the subnet within the supernet.
@@ 2.2 Zero-Cost (ZC) Proxies for NAS
However, the training of the supernet still takes up the majority of the NAS runtime. To speed up the entire NAS process, a more efficient estimation strategy to predict the performance of the searched architecture is needed. Recently, zero-cost proxies have been introduced as a family of performance prediction techniques. The idea is to avoid training any neural network and to run a fast computation (one single forward pass over certain data) over the searched architecture that can assign a ranking or score to each searched architecture.
The expressivity of a neural network, which relates to the complexity of the function it can represent, has been widely used as one of the zero-cost proxies. Several recent works, such as TE-NAS [4], ZEN-NAS[24], and NASWOT [31], approximate the behavior of neural networks by considering the expressivity of ReLU network and the number of linear regions they can separate. Kernel methods have also been employed to approximate the convergence and generalization ability of networks without training. For instance, TE-NAS [4], formulates neural networks as a Gaussian process [17, 52] and analyzes randomly-initialized architectures by the spectrum of the neural tangent kernel (NTK)[11, 48] and the number of linear regions in the input space. Zico[21] extends such kernel-based analysis to reveal the relationships between the gradient properties and the training convergence and generalization capacity of neural networks. Similarly, gradient with respect to the parameters of neural networks is proved to be different approximations of Taylor expansion of deep neural networks [18, 44, 1]. SNIP [18] introduced a saliency metric computed at initialization using a single minibatch of data; Grasp [44] improved upon SNIP [18] by approximating the change in gradient norm instead of loss; Synflow [1] generalized previous approaches and proposed a modified version computes a loss which is simply the product of all parameters in the network thus no data is needed. Gradient concerning feature maps combined with the number of linear regions have been used in Zen-NAS [24]. There are also other zero-cost proxies considering Jacobian covariances, fisher information, and other important theoretical indicators[42, 31, 1]. Although zero-cost proxies can expedite estimation time, they often compromise prediction accuracy and introduce biases stemming from data dependency [46].
2.3 Issue of Model-based Prediction and Unseen Operators
@@
Existing model-based predictor approaches exhibit notable biases, limiting their effectiveness to specific search spaces and requiring fully evaluated architectures as labels. The high initialization time is also a concern. Dudziak et al. [9] attempted to address this issue by leveraging the model’s binary relation and training a prediction model through iterative data selection. However, their approach still requires over 6 GPU days to conduct the search process. Furthermore, due to simplistic architecture and feature embedding methods, such as one-hot encoding, these models cannot process unseen operators. These limitations prompt us to explore a more robust operator encoding method that enables a pretrained prediction model to operate effectively in any architecture search space and accommodate unseen operators. Essentially, we investigate the viability of relying solely on pretrained model-based prediction as a universal zero-cost proxy for guiding searches across diverse architecture spaces with negligible search time.
Figure 1: TG-NAS Framework
3 TG-NAS: A General Zero-Cost Proxy
In this work, we propose TG-NAS, a general zero-cost proxy method for neural architecture search with a transformer embedding generator and a graph neural network trainer, supporting unseen operators in new search spaces. The embedding generator produces the desired feature embeddings with various techniques and the GCN trainer functions as a model-based predictor to predict the accuracy (or ranking) of a given architecture. Finally, the data-independent predicted ranking will guide the search algorithm as shown in Fig. 8.
@@ 3.1 Architecture Representation
Various neural architecture search work represent their networks in the cell structure. For instance, DARTS[28], and NATS-Bench (NAS-Bench-201)[8, 6] define a cell-based search space representing each architecture as a directed acyclic graph (DAG), with nodes representing the features. NAS-Bench-101 [53] on the other hand utilizes nodes to represent layers (operators) and edges for forward dataflow propagation. BANANAS [45] proposed a novel path-based encoding scheme and claimed it scales better than other methods.
In this work, we represent the DNN architecture candidates using DAG with nodes representing operators and edges corresponding to model data propagation flow. We then represent the graphs with adjacent matrix and operator node embeddings, which becomes data input for our graph convolution network (GCN) [14] trainer. To make all search space comply with this representation, TG-NAS unifies other cell-based search space representations as shown in Fig. 8. We applied this transformation to both NAS-Bench-201 and DARTS space in this work. NAS-Bench-101 was provided with the aforementioned architecture graphs, therefore no transformation is needed. Fig. 8 shows an example on NAS-Bench-201. The final model architectures of each search space are obtained by stacking multiple repeated cells with some other predefined cells in between. The differences between different DNN architecture candidates are purely determined by the cell architecture (represented as graph), so we use the embedding of individual cell architectures (as graphs) to represent the entire DNN architecture. Operator node features are encoded using a transformer model with a fixed length embedding size. The details can be seen in the following section.
@@ 3.2 Operator Embedding Generator
Operators Short Sentences Medium Sentences Long Sentences none "None" "Doing nothing" "A none operator that does nothing" skip_connect "Residual connection" "Identity mapping to the next layer" "A residual connection operator that adds identity mapping to the next layer" nor_conv_3x3 "Convolution 3x3" "Convolution 3 by 3 kernel, Batchnorm, ReLU" "A two-dimensional convolutional operator with a kernel size of 3 by 3 is applied, succeeded by a batch normalization layer, and followed by a rectified linear layer" nor_conv_1x1 "Convolution 1x1" "Convolution 1 by 1 kernel, Batchnorm, ReLU" "A two-dimensional convolutional operator with a kernel size of 1 by 1 is applied, succeeded by a batch normalization layer, and followed by a rectified linear layer" avg_pool_3x3 "Average pooling 3x3" "Average pooling 3 by 3 kernel" "A average pooling operator with a kernel size 3 by 3"
@@
Table 2: Results of Different Combination of Sentences Transformer Model and Embedding Sentence Length
@@
Sentence Transformer Model Model Size (MB) Embedding Size Kendall’s $\tau$ Spearman’s $\rho$ Short Medium Long Short Medium Long all-mpnet-base-v2 420 768 0.48 0.49 0.46 0.66 0.67 0.65 MiniLM-L6-v2 80 384 0.57 0.44 0.40 0.77 0.62 0.56 MiniLM-L6-v2-64 81 64 0.36 0.29 0.32 0.54 0.43 0.47
@@
Certain existing works have attempted to construct embedding vectors from words or sentences, such as GloVe [33], or employed character embeddings to capture fine-grained semantic and syntactic regularities. However, our earlier experiments indicated that these methods face challenges when dealing with unseen words, particularly operators in our case. Consequently, we have opted for Sentence BERT [37] as our primary method for generating desired operator embeddings. As illustrated in Fig. 8, the Sentence Transformer utilizes siamese and triplet network structures to derive semantically meaningful sentence embeddings that can be compared using cosine similarity. A pooling operation is applied to the output of the pretrained transformer model to obtain a fixed-sized sentence embedding. We compute the mean of all output word vectors as the pooling strategy to generate the final operator embedding.
@@
In our experiments, we explore three different sizes of pretrained sentence transformer models and three distinct operator sentence lengths, crucial for embedding generation analysis. As outlined in Table 1, we define and experiment with short, medium, and long operator descriptions as inputs to the embedding generator model. The test results on NAS-Bench-201 are presented in Table 2. All three models undergo pretraining on the same 25 dataset collections [37], containing over 1 billion training sentence pairs in total. The triplet objective function is employed during the pretraining phase. This function, given an anchor sentence $a$, a positive sentence $p$, and a negative sentence $n$, tunes the network to ensure that the distance between $a$ and $p$ is smaller than the distance between $a$ and $n$. Mathematically, the loss function can be expressed as:
@@
The all-MiniLM-L6-v2-64 entries in Table 2 are downsampled from the original all-MiniLM-L6-v2 model using principal component analysis (PCA) to achieve a 64-dimensional embedding vector length. The results of our GCN model prediction ranking (which we will discuss in the next section), trained with NAS-Bench-101 and tested on NAS-Bench-201, are presented in Table 2. Notably, the combination of short sentence embedding and the pretrained all-MiniLM-L6-v2-64 sentence transformer model, with embedding length of 384, yields the best results for both Kendall’s $\tau$ and Spearman’s $\rho$ correlation coefficients. Our later experiments will adopt this combination.
Furthermore, we anticipate that fine-tuning the embedding model with context containing specific knowledge about deep learning operators would further enhance the performance of our proposed approach. We leave this area as a potential direction for future research.
3.3 GCN Proxy Trainer

### v2

# TG-NAS: Generalizable Zero-Cost Proxies with Operator Description Embedding and Graph Learning for Efficient Neural Architecture Search

@@ Abstract
Neural Architecture Search (NAS) is a powerful technique for discovering high-performing CNN architectures, but most existing methods rely on costly training or extensive sampling. Zero-shot NAS offers a training-free alternative by using proxies to predict architecture performance. However, existing proxies are often suboptimal—frequently outperformed by simple metrics like parameter count or FLOPs—and they generalize poorly across different search spaces. Moreover, current model-based proxies struggle to adapt to new operators without access to ground-truth accuracy, limiting their transferability. We propose TG-NAS, a universal, model-based zero-cost (ZC) proxy that combines a Transformer-based operator embedding generator with a Graph Convolutional Network (GCN) to predict architecture performance. Unlike prior model-based predictors, TG-NAS requires no retraining and generalizes across arbitrary search spaces. It serves as a standalone ZC proxy with strong data efficiency, robustness, and cross-space consistency. Extensive evaluations across diverse NAS benchmarks demonstrate TG-NAS’s superior rank correlation and generalizability compared to existing proxies. Additional, it improves search efficiency by up to 300× and discovers architectures achieving 93.75% CIFAR-10 accuracy on NAS-Bench-201 and 74.9% ImageNet top-1 accuracy on the DARTS space, establishing TG-NAS as a promising foundation for efficient, generalizable NAS.
@@ 1 Introduction
Deep convolutional neural networks (CNN) have achieved incredible performance in computer vision, speech recognition, object detection, and many other fields [13, 22, 32, 9]. Since then many manually designed CNN topologies that target either high performance or high computational efficiency have been proposed [7, 37]. With larger and deeper expert-designed CNN models, people can easily achieve state-of-the-art performance in various vision tasks. However, manual neural network architecture design is costly and requires a significant amount of time, expertise, and computing resources. It becomes even more challenging if the neural network is designed for hardware platforms with stringent resource and energy constraints. To address the challenges in neural architecture design, neural architecture search (NAS) was proposed to automate the CNN model design flow and explore specialized model architectures given application-specific performance and computing resource constraints [20, 55]. Despite the advancements in the automated search techniques, the majority of neural architecture search (NAS) approaches still face the challenge of time-consuming training and evaluation. In the early stage, prevailing NAS methods typically involved iteratively sampling DNN architectures from the search space, training them, and using their performance to guide the search process [19]. These NAS approaches were primarily based on reinforcement learning (RL) or evolutionary algorithms.
Figure 1: Spearman rank correlation comparison between ours and 4 other ZC proxies (the higher the better).
@@
In recent years, zero-shot proxies [44, 12, 28, 33, 34] have emerged as a promising technique in neural architecture search (NAS). Zero-shot proxies use lightweight computation, typically a single forward pass of the DNN with one minibatch of data, to assign a score to each DNN architecture in the search space. The scores obtained from these proxies are expected to correlate with the final accuracy of candidate DNN architectures [44, 12, 28]. This approach not only conserves computational resources but also accelerates the overall NAS process, enabling researchers to explore a larger number of DNN architectures. However, recent general zero-shot proxies struggle to consistently perform well for practical uses. Many of them cannot outperform a naïve proxy like the number of parameters ($params$) or floating point operation count (FLOPs) as shown in Figure 1, and they are often data-dependent and not fast enough in computation and also do not generalized well across various search space and downstream tasks. Furthermore, the generally high dependence and correlation of different zero-cost (ZC) proxies from similar domains, such as gradient information or kernel methods [14, 4], also reduce the effectiveness of those zero-shot NAS approaches.
@@
To address these limitations of zero-cost performance indicators, many model-based performance predictors have emerged. These predictors rely on training models to forecast the final validation accuracy of an architecture solely based on its topology and operators. Common models include Gaussian processes [16], deep neural networks [39], among others. However, these methods are not zero-shot and cannot be considered zero-cost since they typically require hundreds of fully-trained architectures as training data, resulting in prolonged search times. Additionally, model-based prediction has often been used only to accelerate NAS query times, given their lack of generalization and inability to handle unseen operators that are not encoded during the training of the predictor model.
@@
This leads to fundamental research questions: how can we develop a pure universal prediction model capable of predicting the performance of any given DNN architecture? Is it possible to develop a generally applicable model-based predictor that can handle the unseen operators and mitigate generalization limitations aforementioned? In other words, can we make model-based predictors become a strong standalone zero-cost proxy? Our response to these questions is TG-NAS, a model-based predictor NAS proxy that can be generalized to new search spaces and previously unseen operators without retraining the predictor model. It features a general Transformer-based operator encoder and a graph convolution network (GCN) predictor. TG-NAS provides an efficient and zero-cost (ZC) proxy solution to general neural architecture search problems and bridges the gap between model-based predictors and ZC proxies. We will open source this work to facilitate future research. The link to the source code can be found in the supplementary materials. The main contributions of TG-NAS can be summarized as follows:
@@
  • We propose TG-NAS, a universally applicable, data-independent performance predictor that generalizes to unseen operators and new search spaces without retraining. TG-NAS introduces a novel architecture encoding strategy that integrates: (i) text-based operator descriptions processed by a fine-tuned sentence transformer to capture semantic relationships between operators, (ii) a directed acyclic graph (DAG) based representation of model architectures, and (iii) a GCN-based architecture performance predictor.
  • TG-NAS serves as a zero-cost proxy, enabling prediction-only NAS with no training data dependency. We provide comprehensive analyses spanning 12 NAS benchmarks to demonstrate its superior performance, generalizability and proxy independence.
@@
  • Our analysis of TG-NAS proxy demonstrates remarkable performance on both NAS-Bench-201 and DARTS spaces with CIFAR-10 and ImageNet datasets. TG-NAS achieves up to 300×300$\times$300 × faster search over SOTA zero-cost proxies and magnitudes of speedup over existing NAS methods, while maintaining high accuracy.
@@ 2.1 Neural Architecture Search (NAS)
NAS is an important technique in automating the design of neural architectures for a given task [44]. A typical NAS contains a search strategy that selects the candidate architectures from a predefined search space and an estimation strategy that enables the performance estimation for candidates.
Search space can be categorized as the following types: macro search space [10], chain-structures search space [37], cell-based search space [25, 51, 5], and hierarchical search space [24]. Among those, cell-based search space is the most popular one in NAS [44]. The searchable cells (a directed acyclic graph (DAG) of operations) make up the microstructure of the search space while the macrostructure (that defines the number of cells and how they stack together) is fixed. For example, NAS-Bench-101 [51] contains 423,624 unique architectures, and each cell consists of 7 nodes (each node is chosen from three operators). In NAS-Bench-201 [5], there are 15,625 cell candidates and each cell consists of four nodes (each node chosen from five operators). In contrast, the DARTS search space [25] is more expansive, featuring approximately $10^{18}$ architectures. It consists of two cells, each containing seven nodes. The first two nodes receive inputs from previous layers, and the following four nodes can be any DAG structure, each having two incident edges. The last node serves as the output node, and each edge can take on one of eight operations. In this work, we perform our experiments on those three search spaces to evaluate our proposed TG-NAS flow.
The search strategy in NAS has been widely explored. There are well-known black-box optimizations, such as random selection with full training, reinforcement learning, evolution, Bayesian optimization, Monte Carlo tree search, etc. With these search strategies, we still need to train the searched architecture and use the performance result to guide the search, which is a time-consuming process. To overcome the training time bottleneck, one-shot techniques were introduced as an estimation strategy. These techniques involve training a single (one-shot) supernet, which is an over-parameterized architecture that encompasses all possible architectures in the search space [2, 25]. Once the supernet is trained, each architecture in the search space can be evaluated by inheriting its weights from sampling the subnet within the supernet. Supernet design and training often become the performance bottleneck of these approaches. Supernet training typically dominates the NAS runtime.
@@ 2.2 Zero-Cost (ZC) Proxies for NAS
To speed up the entire NAS process, a more efficient estimation strategy to predict the performance of the searched architecture is needed. Recently, zero-cost proxies have been introduced as a family of performance prediction techniques. The idea is to avoid training any neural network and just run a fast computation, e.g., one single forward pass of the candidate architecture over one tiny set of data, based on which assign a ranking or score to the candidate architecture.
The expressivity of a neural network, which relates to the complexity of the function it can represent, has been widely used as one of the zero-cost proxies. Several recent works, such as TE-NAS [3], Zen-NAS [21], and NASWOT [28], approximate the behavior of neural networks by considering the expressivity of ReLU network and the number of linear regions they can separate. Kernel methods have also been employed to approximate the convergence and generalization ability of networks without training. For instance, TE-NAS [3] formulates neural networks as a Gaussian process and analyzes randomly-initialized architectures by the spectrum of the neural tangent kernel (NTK) [8, 46] and the number of linear regions in the input space. Zico [17] extends such kernel-based analysis to reveal the relationships among the gradient properties, the training convergence, and generalization capacity of neural networks. Similarly, gradient with respect to the parameters of neural networks is proved to be different approximations of Taylor expansion of deep neural networks [14, 42, 1]. SNIP [14] introduced a saliency metric computed at initialization using a single minibatch of data; Grasp [42] improved upon SNIP [14] by approximating the change in gradient norm instead of loss; Synflow [1] generalized previous approaches and proposed a modified version computes a loss which is simply the product of all parameters in the network thus no data is needed. Gradient concerning feature maps combined with the number of linear regions have been used in Zen-NAS [21]. There are also other zero-cost proxies leveraging Jacobian covariances, Fisher information, and other fundamental theoretical indicators [41, 28, 1]. Although zero-cost proxies can expedite estimation time, they often compromise prediction accuracy and introduce biases stemming from data dependency [44].
2.3 Challenges in Model-based Prediction
@@
Existing model-based predictor approaches exhibit notable biases, limiting their effectiveness to specific search spaces and requiring fully evaluated architectures as labels. The high initialization time is also a concern. Dudziak et al. [57] attempted to address this issue by leveraging the model’s binary relation and training a prediction model through iterative data selection. However, their approach still requires over 6 GPU days to conduct the search process. Furthermore, due to simplistic architecture and feature embedding methods, such as one-hot encoding, these models cannot process previously unseen operators. These limitations prompt us to explore a more robust operator encoding method that enables a pre-trained prediction model to operate effectively in any architecture search space and accommodate unseen operators. Essentially, we investigate the viability of relying solely on pre-trained model-based prediction as a universal zero-cost proxy for guiding searches across diverse architecture spaces with negligible search time.
Figure 2: TG-NAS, featuring operator description embedding and GCN-based performance predictor
3 TG-NAS: A Model-based Zero-cost Proxy
In this work, we propose TG-NAS, a general model-based zero-cost proxy method for neural architecture search with a transformer embedding generator and a graph neural network predictor, supporting unseen operators in new search spaces. The embedding generator produces the desired feature embeddings with various techniques and the GCN predictor was trained with existing NAS benchmarks to predict the accuracy (or ranking) of a given architecture. Finally, the data-independent predicted ranking will guide the search algorithm as shown in Figure 2.
@@ 3.1 Architecture Representation
Various neural architecture search works represent their networks in the cell structure. For instance, DARTS [25], and NATS-Bench (NAS-Bench-201) [5] define a cell-based search space representing each architecture as a directed acyclic graph (DAG), with nodes representing the features. NAS-Bench-101 [51] on the other hand utilizes nodes to represent layers (operators) and edges for forward dataflow propagation. BANANAS [43] proposed a novel path-based encoding scheme and claimed it scales better than other methods. Additionally, Yan et al.  [50] propose a transformer-based encoding scheme with computation awareness. On contrary, our transformer-based neural architecture coding uses text-based DNN operator descriptions and sentence transformer.
In this work, we represent the DNN architecture candidates using DAG with nodes representing operators and edges corresponding to model data propagation flow. We then represent the graphs with adjacent matrix and operator node embeddings, which becomes data input for our graph convolution network (GCN)[11] predictor. To make all search space comply with this representation, TG-NAS unifies other cell-based search space representations as shown in Figure 2. We applied this transformation to all NAS benchmark and search space in this work. NAS-Bench-101 was provided with the aforementioned architecture graphs, therefore no transformation is needed. Figure 2 shows an example on NAS-Bench-201. The final model architectures of each search space are obtained by stacking multiple repeated cells with some other predefined cells in between. The differences between different DNN architecture candidates are purely determined by the cell architecture (represented as graph), so we use the embedding of individual cell architectures (as graphs) to represent the entire DNN architecture. Operator node features are encoded using a transformer model with a fixed length embedding size. The details can be seen in the following section.
@@ 3.2 Operator Embedding Generator
Operators Short Sentences Medium-length Sentences Long Sentences none “None” “Doing nothing” “A none operator that does nothing” skip_connect “Residual connection" “Identity mapping to the next layer" “A residual connection operator that adds identity mapping to the next layer" nor_conv_3x3 “Convolution 3x3" “Convolution 3 by 3 kernel, Batchnorm, ReLU" “A two-dimensional convolutional operator with a kernel size of 3 by 3 is applied, succeeded by a batch normalization layer, and followed by a rectified linear layer" nor_conv_1x1 “Convolution 1x1" “Convolution 1 by 1 kernel, Batchnorm, ReLU" “A two-dimensional convolutional operator with a kernel size of 1 by 1 is applied, succeeded by a batch normalization layer, and followed by a rectified linear layer" avg_pool_3x3 “Average pooling 3x3" “Average pooling 3 by 3 kernel" “A average pooling operator with a kernel size 3 by 3"
@@
Table 2: Different Combination of Sentences Transformer Model and Embedding Sentence Length
@@
Model Model Size (MB) Embedding Size Kendall’s $\tau$ Spearman’s $\rho$ Short Medium Long Short Medium Long all-mpnet-base-v2 420 768 0.48 0.49 0.46 0.66 0.67 0.65 MiniLM-L6-v2 80 384 0.56(0.60*) 0.44 0.40 0.76(0.80*) 0.62 0.56 MiniLM-L6-v2-64 81 64 0.36 0.29 0.32 0.54 0.43 0.47
*Fine Tuned with Augmented DNN Operator Descriptions
@@
Certain existing works have attempted to construct embedding vectors from words or sentences, such as GloVe [30], or employed character embeddings to capture fine-grained semantic and syntactic regularities. However, our earlier experiments indicated that these methods face challenges when dealing with previously unseen words, particularly operators in our case. Consequently, we have opted for Sentence Transformer [36] as our primary method for generating desired operator embeddings. As illustrated in Figure 2, the Sentence Transformer utilizes siamese and triplet network structures to derive semantically meaningful sentence embeddings that can be compared using cosine similarity. A pooling operation is applied to the output of the pre-trained transformer model to obtain a fixed-size sentence embedding. We compute the mean of all output word vectors as the pooling strategy to generate the final operator embedding.
@@
In our experiments, we explore three different sizes of pre-trained sentence transformer models and three distinct operator sentence lengths, crucial for embedding generation analysis. As outlined in Table 2, we define and experiment with short, medium, and long operator descriptions as inputs to the embedding generator model. The test results on NAS-Bench-201 are presented in Table 2. All three models undergo pretraining on the same 25 dataset collections [36], containing over 1 billion training sentence pairs in total. The triplet objective function is employed during the pretraining phase. This function, given an anchor sentence $a$, a positive sentence $p$, and a negative sentence $n$, tunes the network to ensure that the distance between $a$ and $p$ is smaller than the distance between $a$ and $n$. Mathematically, the loss function can be expressed as:
@@
The all-MiniLM-L6-v2-64 entries in Table 2 are downsampled from the original all-MiniLM-L6-v2 model using principal component analysis (PCA) to achieve a 64-dimensional embedding vector length. The results of our GCN model prediction ranking (which we will discuss in the next section), trained with NAS-Bench-101 and tested on NAS-Bench-201, are presented in Table 2. Notably, the combination of short sentence embedding and the pre-trained all-MiniLM-L6-v2 sentence transformer model, with the embedding length of 384, yields the best results for both Kendall’s $\tau$ and Spearman’s $\rho$ correlation coefficients. Our later experiments will adopt this combination.
Furthermore, to tailor the embedding space more closely to neural architecture search, we fine-tune the MiniLM-L6-v2 model on a task-specific similarity dataset, which is constructed from the PyTorch torch.nn documentation. Specifically, we parse operator classes and group them by functional categories (e.g., convolution, pooling, normalization, activation). For each operator class, we extract both the class name and its accompanying textual description. To enhance both the diversity and semantic richness of the dataset, we further employ GPT-4o to generate multiple augmented descriptions per class. These augmentations provide varied linguistic formulations that better reflect how operator functionality can be described in natural language, thus helping the model generalize across heterogeneous textual inputs during inference. We fine-tune the operator embedding model using supervised similarity pairs constructed based on functional categories. Pairs from the same operator class are labeled as similar, pairs from different classes within the same category as related, and pairs from distinct categories as unrelated. A cosine similarity loss is used to align the embedding space accordingly. Detailed definitions are provided in Appendix.
Algorithm 1 TG-NAS Search Algorithm
1:Input: $NP$ population size
2:$g\leftarrow 0$
3:while $|pop|<NP$ do
[…]
5:    $pop^{\prime}_{i}\leftarrow\text{discretized\_architecture}(pop_{i})$
6:    $fitness_{i}\leftarrow\text{GCN\_Inference}(pop^{\prime}_{i})$
7:end while
8:while $g<g_{max}$ do
9:    $V_{g}\leftarrow\text{mutate}(pop_{g})$
10:    $U_{g}\leftarrow\text{crossover}(V_{g},pop_{g})$
11:    $U^{\prime}_{g}\leftarrow\text{discretized\_population}(U_{g})$
12:    $fitness_{g}\leftarrow\text{GCN\_Inference}(U^{\prime}_{g})$
13:    $pop_{g+1},fitness_{g+1}\leftarrow\text{select}(pop_{g},U_{g})$
14:end while
15:return $POP_{g}$
3.3 GCN Proxy Predictor